### v1

@@ 2.3 Scaling relations
The anticorrelation between optical variability and luminosity, in the sense that most luminous quasars are less variable than fainter ones, was discovered very early on [24]. This anticorrelation was later confirmed by studies based on tens to hundreds of quasars [e.g. 55, 110, 291, 53, 54, 95], optical [125, 256, 58] and X-ray selected AGN [279], as well as ensemble studies of thousands of quasars from the QUEST1, SDSS, Palomar-Quest, Pan-STARRS1, DECaLS, CRTS and PTF surveys [e.g. 314, 59, 316, 315, 23, 190, 329, 177, 145, 47, 156, 282, 163, 241]. The anticorrelation can be interpreted as a consequence of the fact that the most luminous AGN are likely to host larger mass black holes (if all quasars accrete at similar Eddington rates). In this case, their variability should be diluted by the larger emitting region.
@@
Regarding the accretion rate, most works suggest the presence of an anticorrelation between the variability amplitude and the accretion rate [e.g. 23, 21, 190, 279, 315, 256, 156, 163, 58, 322, 6, 241], but it is often the case that this result is degenerate with the luminosity dependence.
@@
The above scaling relations generally apply to the variability amplitude, but when the available data allow to constrain them, the characteristic variability timescales also appear to be correlated with AGN physical parameters. For example [48, 125, 190, 145, 156, 273, 282, 38, 37, 268, 6, 288] report a positive correlation with MBH and/or with , derived from DRW/DHO fitting, SF or PSD analysis (Fig. 10). Graham et al. [89], Yu et al. [322] also observe an anticorrelation with bolometric luminosity, while other works [e.g. 21, 279] do not observe a significant correlation between the characteristic timescale and the other parameters of the system. Notice, however, that what is often claimed to be a characteristic timescale is actually defined at fixed amplitude, thus preventing a direct and meaningful comparison between various works. Kozłowski [145] also found a mild dependence of the SF slope on luminosity and a lack of dependence on MBH, while Li et al. [156] find a positive correlation with both. We caution, however, that such results may depend on the specific approach used to model the light curves (e.g. DRW, DHO, etc.) or on the maximum temporal baseline probed by each sample, as discussed previously. A summary of some of these contrasting results is reported in Suberlak et al. [268] and De Cicco et al. [58], and is revised and expanded here in Table 1.
@@
Clearly, much of the confusion arises from the use of different variability estimators, different samples, and the inability to remove the degeneracies among physical properties, intrinsic or observational, due to the limited parameter space accessible to most studies so far. For instance, the SDSS Stripe 82 study of [241, see also 190, 322 and references therein] clearly shows how the availability of long lightcurves in multiple bands and the relatively large range of redshifts, masses and luminosities of the quasars in the sample, allows one to properly account for all such effects. Coupled with a homogeneous analysis of the quasar power spectra, it can reveal clear trends in both mass and accretion rate (or luminosity), as shown in Fig. 11. Most notably, the dependence on BH mass and accretion rate appear to be independent from each other, although, strictly speaking, this result is limited to the range of temporal frequencies sampled by the Stripe 82 sample.
@@
These results have led several authors to speculate that the observed variability properties of quasars (and possibly also lower luminosity AGN) can be explained if there is a universal PSD shape in all AGN and that the characteristic timescales depend on the physical properties of the source (i.e. BH mass and accretion rate). Tang et al. [295, also see ] suggest that the quasar SF in the UV band from the ATLAS survey is consistent with a universal random walk if time is expressed in units of the orbital or thermal timescales of the emitting material. On the other hand, [6] apply the scaling relations discussed above to build a universal power spectrum of ZTF quasars which can be modelled by a bending power-law and appears steeper than DRW predictions, showing no evidence of flattening at the lowest probed frequencies (Fig. 12, left). [241] also proposed that the ensemble PSD of SDSS quasars is consistent with a universal bending power-law shape, when frequencies are rescaled using the gravitational timescale (Fig. 12, right).
@@
Works based on non-parametric approaches using ML algorithms (Sect. 2.1) also suggest that additional information about the physical properties of the accreting system may be embedded in the detailed shape of the lightcurves themselves, instead of in the global variability amplitude and timescale. Asymmetries in the brightening and fading phases may in fact depend on the AGN mass and luminosity [285, and references therein].
@@ 2.4 Interpretation of AGN optical variability
where is the radial distance and is the mass of the central BH. This timescale is small ( a few days) even for the largest SMBH and if we consider the disc regions up to RS. According to the standard Shakura-Sunyaev [276] accretion disk model (which is still debated for AGN), the temperature profile of a geometrically thin Keplerian disc scales as . The disc emission in a particular waveband will originate from a broad range of radii, and could be affected by various time-scales that determine the flux variations from accretion discs [see, e.g. 84]. For example, the emitted flux can be modulated on the disc orbital timescale:
@@
where is the viscosity of the disk. Thus for a fixed viscosity, the thermal timescale is proportional to the orbital timescale. For typical BHs and then days-years at . These values appear to be consistent claims of breaks (i.e. characteritic timescales) detected in the PSD/SF of AGN [e.g. 48, 273].
Works based on DRW lightcurve fitting or SF analysis also propose that the measured characteristic timescales of quasars are in agreement with the disc dynamical or thermal timescales, assuming a viscosity parameter of [e.g. 125, 37, 295]. The thermal hypothesis is further supported by the fact that there are no clear periodicities, as could be expected from Keplerian motions, and by the trend of the observed emission to become bluer when brighter (Sect. 2.2). In any case, both timescales intrinsically include a dependence on the mass, which may be imprinted in the correlations discussed in Sect. 2.3. However, they do not depend on the accretion rate of the disc thus failing to explain in full the observed dependencies.
@@
Thus the viscous and sound timescales may explain the dependence of the characteristic timescales on accretion rate, as it is claimed by several studies. In general if of an object is close to 1, remains relatively small above but for super-Eddington flows, or for a hot, optically-thin plasma at the virial temperature this ratio is close to 1, and the viscous timescale is equal to the thermal timescale. There are additional possible timescales allowed by current accretion models, such as the “Cold Disk Removal” timescale, due to the radius where the disk changes from optically-thick to optically-thin [56], the free-fall timescale that may characterise Advection-Dominated Accretion flows [ADAF, 273].
@@
While these basic physical considerations provide a useful framework for understanding AGN variability, it is likely that the actual scenario is more complicated, and 3D Magneto-Hydrodynamical (MHD) simulations are needed to fully understand the disk structure. Although this is a complex topic that we cannot discuss here in detail (see Davis and Tchekhovskoy 70 for a recent review), we point out that Magneto-Rotational Instability (MRI) is believed to play a major role in generating the disk viscosity and in producing local fluctuations in dissipation and accretion rate, as well as disk clumpiness and vertical stratification. MHD simulations predict that the formation of large magnetic loops takes more than one dynamical timescale, so that the magnetic field affects the X-ray corona and the disk evolution on timescales between the dynamical and thermal one [e.g. 200]. More recent work also shows that when the metal opacity is properly accounted for, the disk may become convectively unstable. This produces strong fluctuations in surface density and disk heating resulting in significant luminosity variations, which may explain the observed AGN variability over a timescale of years to decades (Fig. 13, Jiang and Blaes 120, and references therein).
@@
In any case, since the observed emission at every wavelength is due to a superposition of contributions from different parts of the disk, in reality it is questionable if we should expect to observe a dominant timescale in AGN lightcurves (although we do expect the SF/PSD to flatten on long timescales/low frequencies to prevent a diverging total power). In fact, despite several claims (Sect. 2.2), current observations still did not definitively prove the presence of a unique timescale probing the physical properties of the disk (differently from X-rays, see Sect. 3) while the case for a correlation of the variability amplitude with physical parameters appears stronger. Clearly a superposition of different effects cannot be excluded and is actually favoured by many studies. In particular, the plasma may respond to non-local perturbations. These include irradiation of the outer disk by radiation generated in the inner parts of the disk itself or the optically thin plasma in the central regions (the “corona”), or mechanical transport of energy by convection or wind/coronal outflows [see 56, and references therein].
@@
Microlensing effects caused by intervening bodies within the host galaxy or in the galactic halo may also contribute to the observed variability on short timescales. This form of variability enables mapping of quasar structure down to arcsec scales and can be best constrained in lensed quasars, and has suggested that the accretion disk sizes are larger than expected for a thin-disc, requiring to consider extensions/modifications of the standard Shakura-Sunyaev model or the contribution of additional diffuse emission originating from, e.g., the Broad-Line Region (313 for a recent review on the topic). However when proper relativistic effects are taken into account, as further discussed below in the context of X-ray reverberation, this tension can be reconciled with the standard disk model.
@@
In fact, in the last decade, there have been many multi-wavelength monitoring campaigns with the goal of studying the correlation between X-ray, UV, and optical variations in AGN. The targets are mainly near bright Seyferts, which are monitored mainly by the Neil Gherels Swift Observatory and AstroSat, NICER (to a lesser extent) as well as ground-based telescopes in some cases [e.g. 180, 75, 74, 73, 82, 43, 45, 107, 311, 141, 124, 62, 181, 131, 132, 225]. Figure 3 shows the resulting light curves in the case of the NGC 5548 STORM 1, multi-wavelength monitoring campaign. Similar quality light curves (in terms of sampling rate and duration) have resulted from most of the other campaigns as well. In all cases, the UV/optical variations are well correlated, and the variations propagate from the shorter wavelength to the longer wavelength. Well-correlated variations propagating from the UV to the optical/IR bands are also observed in AGN which are at higher redshifts and brighter than their nearby counterparts [e.g. 122, 195, 114, 113, 321, 88, 123]. This is an important observational result that puts strong constraints on theoretical models. Any model for the UV/optical variability of AGN should be able to explain these correlations which appear to hold whenever variable AGN are monitored in many wavebands.
@@
These results are consistent with the hypothesis of X-ray reverberation in AGN. According to this model, part of the X-rays emitted by the corona in AGN are directed towards the accretion disc. A fraction is ‘reflected’ by the disc in the X-ray range (this is the so-called “X-ray reflection component” in AGN). The rest is absorbed, increasing the temperature of the disc and consequently the emitted UV/optical flux as well. Since X-rays are highly variable, the additional UV/optical emission of the disc will also be variable, and the optical/UV variations will be delayed with respect to the X-rays. Since the disc emitting area increases with increasing wavelength, we expect the time lags to also increase with wavelength (as observed). This phenomenon is referred to as the “disc X-ray thermal reverberation” [e.g. 142, 46].
@@
In the case of a standard [276] accretion disc, we expect the X-ray/UV/optical time delays, , to be wavelength dependent following a relation of the form [see 46]. The cross-correlation results from the first monitoring campaigns confirmed this prediction. However, it appeared that the amplitude of this relation underestimated the observed lags by a factor of 3-4 (see, e.g., the left panel of Fig. 4). [130], however, studied X-ray reverberation in detail and showed that this is not the case [see also 69]. They assumed an X-ray point-like source that illuminates a standard Novikov-Thorne accretion disc [212]. They considered in detail all special and general relativity effects in the light propagation from the X-ray corona to the disc and from the disc to the observer, and they took into account the disc ionisation state when computing the disc reflection flux. As a result, they were able to compute the exact amount of X-rays that are absorbed by the disc in each radius. This was a major advantage compared to previous works, which simply assumed equal power for the X-ray absorbed flux and the internally generated heat in the disc at each radius. Using the [130] relations, [147] showed that the observed time lags in the nearby Seyferts are fully consistent with X-ray reverberation (see, for example, the right panel in Fig. 4 for the case of NGC 5548). The same is true for the observed time lags of the more distant AGN [158].
@@
[150] provided a code that can compute X-ray reverberation time lags for a large parameter space and fitted the observed optical/UV time lags in many Seyferts (left panel, Fig. 14). X-ray reverberation models can also fit the wavelength dependence of the observed variance in nearby AGN [240], the apparent lack of significant cross-correlation between X-ray and optical/UV [226], the frequency-dependent time lags in the optical/UV band [239], the optical/UV spectral variability observed in NGC 5548 [148], and the accretion disc size problems discussed above, inferred from microlensing observations [221]. The hypothesis of X-ray disc thermal reverberation can also fit well the optical/UV power spectra of NGC 5548 [238] (right panel, Fig. 14). At least in this source, (almost) all of the observed UV and optical variations could be due to X-rays illuminating the disc. This possibility alleviates (at least in this source) the “problem” of the fast optical/UV variations that are observed in many AGN, and cannot be explained by the standard accretion disc timescales.
@@
Optical/UV time lags similar to what we observe can be produced by line and continuum emission from gas in the BLR which responds to the continuum extrem UV variations [e.g. 133, 208]. In fact, [208] shows that the optical/UV time-lag spectra for many AGN are consistent with the response of diffuse emission from radiation pressure-supported clouds in the BLR with a covering factor of about 0.2, without any need for disc X-ray thermal reverberation (left panel of Fig. 15). However, this is not the case with the Fairall 9 time-lags [81]. A complete lack of disc X-ray reverberation would be achieved only if the X-ray source is completely shielded from the disc, by an unknown structure, or in a geometry where the X-ray corona is located within the disc and its height is very small. In this case, the extreme UV part of the disc must be variable to be able to drive the variations of the BLR clouds.

### v3

@@ 2.3 Scaling relations
The anticorrelation between optical variability and luminosity, in the sense that most luminous quasars are less variable than fainter ones, was discovered very early on [25]. This anticorrelation was later confirmed by studies based on tens to hundreds of quasars [e.g. 56, 112, 293, 54, 55, 97], optical [127, 258, 59] and X-ray selected AGN [281], as well as ensemble studies of thousands of quasars from the QUEST1, SDSS, Palomar-Quest, Pan-STARRS1, DECaLS, CRTS and PTF surveys [e.g. 316, 60, 318, 317, 24, 192, 331, 179, 147, 48, 158, 284, 165, 243]. The anticorrelation can be interpreted as a consequence of the fact that the most luminous AGN are likely to host larger mass black holes (if all quasars accrete at similar Eddington rates). In this case, their variability should be diluted by the larger emitting region.
@@
Regarding the accretion rate, most works suggest the presence of an anticorrelation between the variability amplitude and the accretion rate [e.g. 24, 22, 192, 281, 317, 258, 158, 165, 59, 324, 6, 243], but it is often the case that this result is degenerate with the luminosity dependence.
@@
The above scaling relations generally apply to the variability amplitude, but when the available data allow to constrain them, the characteristic variability timescales also appear to be correlated with AGN physical parameters. For example [49, 127, 192, 147, 158, 275, 284, 39, 38, 270, 6, 290] report a positive correlation with MBH and/or with , derived from DRW/DHO fitting, SF or PSD analysis (Fig. 10). Graham et al. [91], Yu et al. [324] also observe an anticorrelation with bolometric luminosity, while other works [e.g. 22, 281] do not observe a significant correlation between the characteristic timescale and the other parameters of the system. Notice, however, that what is often claimed to be a characteristic timescale is actually defined at fixed amplitude, thus preventing a direct and meaningful comparison between various works. Kozłowski [147] also found a mild dependence of the SF slope on luminosity and a lack of dependence on MBH, while Li et al. [158] find a positive correlation with both. We caution, however, that such results may depend on the specific approach used to model the light curves (e.g. DRW, DHO, etc.) or on the maximum temporal baseline probed by each sample, as discussed previously. A summary of some of these contrasting results is reported in Suberlak et al. [270] and De Cicco et al. [59], and is revised and expanded here in Table 1.
@@
Clearly, much of the confusion arises from the use of different variability estimators, different samples, and the inability to remove the degeneracies among physical properties, intrinsic or observational, due to the limited parameter space accessible to most studies so far. For instance, the SDSS Stripe 82 study of [243, see also 192, 324 and references therein] clearly shows how the availability of long lightcurves in multiple bands and the relatively large range of redshifts, masses and luminosities of the quasars in the sample, allows one to properly account for all such effects. Coupled with a homogeneous analysis of the quasar power spectra, it can reveal clear trends in both mass and accretion rate (or luminosity), as shown in Fig. 11. Most notably, the dependence on BH mass and accretion rate appear to be independent from each other, although, strictly speaking, this result is limited to the range of temporal frequencies sampled by the Stripe 82 sample.
@@
These results have led several authors to speculate that the observed variability properties of quasars (and possibly also lower luminosity AGN) can be explained if there is a universal PSD shape in all AGN and that the characteristic timescales depend on the physical properties of the source (i.e. BH mass and accretion rate). Tang et al. [297, also see ] suggest that the quasar SF in the UV band from the ATLAS survey is consistent with a universal random walk if time is expressed in units of the orbital or thermal timescales of the emitting material. On the other hand, [6] used the Mexican-hat method of [7] to compute power spectra of quasars using ZTF light curves. Combined with the scaling relations discussed above, they built a universal power spectrum which can be modelled by a bending power law and appears steeper than DRW predictions, showing no evidence of flattening at the lowest probed frequencies (Fig.,12, left). [243] also proposed that the ensemble PSD of SDSS quasars is consistent with a universal bending power-law shape, when frequencies are rescaled using the gravitational timescale (Fig. 12, right).
@@
Works based on non-parametric approaches using ML algorithms (Sect. 2.1) also suggest that additional information about the physical properties of the accreting system may be embedded in the detailed shape of the lightcurves themselves, instead of in the global variability amplitude and timescale. Asymmetries in the brightening and fading phases may in fact depend on the AGN mass and luminosity [287, and references therein].
@@ 2.4 Interpretation of AGN optical variability
where is the radial distance and is the mass of the central BH. This timescale is small ( a few days) even for the largest SMBH and if we consider the disc regions up to RS. According to the standard Shakura-Sunyaev [278] accretion disk model (which is still debated for AGN), the temperature profile of a geometrically thin Keplerian disc scales as . The disc emission in a particular waveband will originate from a broad range of radii, and could be affected by various time-scales that determine the flux variations from accretion discs [see, e.g. 86]. For example, the emitted flux can be modulated on the disc orbital timescale:
@@
where is the viscosity of the disk. Thus for a fixed viscosity, the thermal timescale is proportional to the orbital timescale. For typical BHs and then days-years at . These values appear to be consistent claims of breaks (i.e. characteritic timescales) detected in the PSD/SF of AGN [e.g. 49, 275].
Works based on DRW lightcurve fitting or SF analysis also propose that the measured characteristic timescales of quasars are in agreement with the disc dynamical or thermal timescales, assuming a viscosity parameter of [e.g. 127, 38, 297]. The thermal hypothesis is further supported by the fact that there are no clear periodicities, as could be expected from Keplerian motions, and by the trend of the observed emission to become bluer when brighter (Sect. 2.2). In any case, both timescales intrinsically include a dependence on the mass, which may be imprinted in the correlations discussed in Sect. 2.3. However, they do not depend on the accretion rate of the disc thus failing to explain in full the observed dependencies.
@@
Thus the viscous and sound timescales may explain the dependence of the characteristic timescales on accretion rate, as it is claimed by several studies. In general if of an object is close to 1, remains relatively small above but for super-Eddington flows, or for a hot, optically-thin plasma at the virial temperature this ratio is close to 1, and the viscous timescale is equal to the thermal timescale. There are additional possible timescales allowed by current accretion models, such as the “Cold Disk Removal” timescale, due to the radius where the disk changes from optically-thick to optically-thin [57], the free-fall timescale that may characterise Advection-Dominated Accretion flows [ADAF, 275].
@@
While these basic physical considerations provide a useful framework for understanding AGN variability, it is likely that the actual scenario is more complicated, and 3D Magneto-Hydrodynamical (MHD) simulations are needed to fully understand the disk structure. Although this is a complex topic that we cannot discuss here in detail (see Davis and Tchekhovskoy 72 for a recent review), we point out that Magneto-Rotational Instability (MRI) is believed to play a major role in generating the disk viscosity and in producing local fluctuations in dissipation and accretion rate, as well as disk clumpiness and vertical stratification. MHD simulations predict that the formation of large magnetic loops takes more than one dynamical timescale, so that the magnetic field affects the X-ray corona and the disk evolution on timescales between the dynamical and thermal one [e.g. 202]. More recent work also shows that when the metal opacity is properly accounted for, the disk may become convectively unstable. This produces strong fluctuations in surface density and disk heating resulting in significant luminosity variations, which may explain the observed AGN variability over a timescale of years to decades (Fig. 13, Jiang and Blaes 122, and references therein).
@@
In any case, since the observed emission at every wavelength is due to a superposition of contributions from different parts of the disk, in reality it is questionable if we should expect to observe a dominant timescale in AGN lightcurves (although we do expect the SF/PSD to flatten on long timescales/low frequencies to prevent a diverging total power). In fact, despite several claims (Sect. 2.2), current observations still did not definitively prove the presence of a unique timescale probing the physical properties of the disk (differently from X-rays, see Sect. 3) while the case for a correlation of the variability amplitude with physical parameters appears stronger. Clearly a superposition of different effects cannot be excluded and is actually favoured by many studies. In particular, the plasma may respond to non-local perturbations. These include irradiation of the outer disk by radiation generated in the inner parts of the disk itself or the optically thin plasma in the central regions (the “corona”), or mechanical transport of energy by convection or wind/coronal outflows [see 57, and references therein].
@@
Microlensing effects caused by intervening bodies within the host galaxy or in the galactic halo may also contribute to the observed variability on short timescales. This form of variability enables mapping of quasar structure down to arcsec scales and can be best constrained in lensed quasars, and has suggested that the accretion disk sizes are larger than expected for a thin-disc, requiring to consider extensions/modifications of the standard Shakura-Sunyaev model or the contribution of additional diffuse emission originating from, e.g., the Broad-Line Region (315 for a recent review on the topic). However when proper relativistic effects are taken into account, as further discussed below in the context of X-ray reverberation, this tension can be reconciled with the standard disk model.
@@
In fact, in the last decade, there have been many multi-wavelength monitoring campaigns with the goal of studying the correlation between X-ray, UV, and optical variations in AGN. The targets are mainly near bright Seyferts, which are monitored mainly by the Neil Gherels Swift Observatory, AstroSat and NICER (to a lesser extent) as well as ground-based telescopes in some cases [e.g. 182, 77, 76, 75, 84, 44, 46, 109, 313, 143, 126, 63, 183, 133, 134, 227]. Figure 3 shows the resulting light curves in the case of the NGC 5548 STORM 1, multi-wavelength monitoring campaign. Similar quality light curves (in terms of sampling rate and duration) have resulted from most of the other campaigns as well. In all cases, the UV/optical variations are well correlated, and the variations propagate from the shorter wavelength to the longer wavelength. Well-correlated variations propagating from the UV to the optical/IR bands are also observed in AGN which are at higher redshifts and brighter than their nearby counterparts [e.g. 124, 197, 116, 115, 323, 90, 125]. This is an important observational result that puts strong constraints on theoretical models. Any model for the UV/optical variability of AGN should be able to explain these correlations which appear to hold whenever variable AGN are monitored in many wavebands.
@@
These results are consistent with the hypothesis of X-ray reverberation in AGN. According to this model, part of the X-rays emitted by the corona in AGN are directed towards the accretion disc. A fraction is ‘reflected’ by the disc in the X-ray range (this is the so-called “X-ray reflection component” in AGN). The rest is absorbed, increasing the temperature of the disc and consequently the emitted UV/optical flux as well. Since X-rays are highly variable, the additional UV/optical emission of the disc will also be variable, and the optical/UV variations will be delayed with respect to the X-rays. Since the disc emitting area increases with increasing wavelength, we expect the time lags to also increase with wavelength (as observed). This phenomenon is referred to as the “disc X-ray thermal reverberation” [e.g. 144, 47].
@@
In the case of a standard [278] accretion disc, we expect the X-ray/UV/optical time delays, , to be wavelength dependent following a relation of the form [see 47]. The cross-correlation results from the first monitoring campaigns confirmed this prediction. However, it appeared that the amplitude of this relation underestimated the observed lags by a factor of 3-4 (see, e.g., the left panel of Fig. 4). [132], however, studied X-ray reverberation in detail and showed that this is not the case [see also 71]. They assumed an X-ray point-like source that illuminates a standard Novikov-Thorne accretion disc [214]. They considered in detail all special and general relativity effects in the light propagation from the X-ray corona to the disc and from the disc to the observer, and they took into account the disc ionisation state when computing the disc reflection flux. As a result, they were able to compute the exact amount of X-rays that are absorbed by the disc in each radius. This was a major advantage compared to previous works, which simply assumed equal power for the X-ray absorbed flux and the internally generated heat in the disc at each radius. Using the [132] relations, [149] showed that the observed time lags in the nearby Seyferts are fully consistent with X-ray reverberation (see, for example, the right panel in Fig. 4 for the case of NGC 5548). The same is true for the observed time lags of the more distant AGN [160].
@@
[152] provided a code that can compute X-ray reverberation time lags for a large parameter space and fitted the observed optical/UV time lags in many Seyferts (left panel, Fig. 14). X-ray reverberation models can also fit the wavelength dependence of the observed variance in nearby AGN [242], the apparent lack of significant cross-correlation between X-ray and optical/UV [228], the frequency-dependent time lags in the optical/UV band [241], the optical/UV spectral variability observed in NGC 5548 [150], and the accretion disc size problems discussed above, inferred from microlensing observations [223]. The hypothesis of X-ray disc thermal reverberation can also fit well the optical/UV power spectra of NGC 5548 [240] (right panel, Fig. 14). At least in this source, (almost) all of the observed UV and optical variations could be due to X-rays illuminating the disc. This possibility alleviates the “problem” of the fast optical/UV variations that are observed in many AGN, and cannot be explained by the standard accretion disc timescales.
@@
Optical/UV time lags similar to what we observe can be produced by line and continuum emission from gas in the BLR which responds to the continuum extrem UV variations [e.g. 135, 210]. In fact, [210] shows that the optical/UV time-lag spectra for many AGN are consistent with the response of diffuse emission from radiation pressure-supported clouds in the BLR with a covering factor of about 0.2, without any need for disc X-ray thermal reverberation (left panel of Fig. 15). However, this is not the case with the Fairall 9 time-lags [83]. A complete lack of disc X-ray reverberation would be achieved only if the X-ray source is completely shielded from the disc, by an unknown structure, or in a geometry where the X-ray corona is located within the disc and its height is very small. In this case, the extreme UV part of the disc must be variable to be able to drive the variations of the BLR clouds.